Suppose the risk-free rate is 2.40% and an analyst assumes a market risk premium of 6.68%. Firm A just paid a dividend of $1.12 per share. The analyst estimates the β of Firm A to be 1.35 and estimates the dividend growth rate to be 4.91% forever. Firm A has 264.00 million shares outstanding. Firm B just paid a dividend of $1.53 per share. The analyst estimates the β of Firm B to be 0.71 and believes that dividends will grow at 2.91% forever. Firm B has 188.00 million shares outstanding. What is the value of Firm B? (ANSWER SHOULD BE IN THE MILLIONS or BILLIONS.)
Answer #1

Value of Firm B=Number of shares*D0*(1+g)/(rf+b*MRP-g)=188*1.53*1.0291/(2.40%+0.71*6.68%-2.91%)=6993.250898 million